OpenStudy (steve816):

Please help I'm really struggling! How many miles/hour is 30 km/s?

OpenStudy (anonymous):

1 km =0.625 miles 1s=1/3600 hour

OpenStudy (steve816):

I got 67,068 mile/hour, can someone check my answer?

jimthompson5910 (jim_thompson5910):

1 km = 0.621371 mi (approx) 1 hr = 3600 seconds \[\Large \frac{30 \text{ km}}{1 \text{ sec}} \times \frac{0.621371 \text{ mi}}{1 \text{ km}} \times \frac{3600 \text{ sec}}{1 \text{ hr}}\] Notice how I set this up to where the "km" units cancel (one "km" unit up top, the other "km" down in the denominator. They divide to 1) \[\Large \frac{30 \cancel{\text{ km}}}{1 \text{ sec}} \times \frac{0.621371 \text{ mi}}{1 \cancel{\text{ km}}} \times \frac{3600 \text{ sec}}{1 \text{ hr}}\] Also, the "sec" units cancel as well for the same reason as the "km" units cancelling \[\Large \frac{30 \text{ km}}{1 \cancel{\text{ sec}}} \times \frac{0.621371 \text{ mi}}{1 \text{ km}} \times \frac{3600 \cancel{\text{ sec}}}{1 \text{ hr}}\]

jimthompson5910 (jim_thompson5910):

`I got 67,068 mile/hour, can someone check my answer?` I'm getting a slightly larger number @steve816

OpenStudy (steve816):

I think it's because I rounded the miles to 0.62. But I understand it now! Thanks for the help!

jimthompson5910 (jim_thompson5910):

no problem